# On the Roles of LLMs in Planning: Embedding LLMs into Planning Graphs

### v1

@@ 1 Introduction
Plan synthesis aims to generate a course of actions or policies to transit given initial states to goal states, provided domain models that could be designed by experts or learnt from training data (Aineto et al., 2019) or interactions with the world (Lamanna et al., 2023; Jin et al., 2022b). It is a time- and space-consuming open issue in the planning community (Ghallab et al., 2004). Intrigued by the claims of emergent planning capabilities in large language models (LLMs), works have been proposed to investigate the planning effectiveness of LLMs, without considering any utilization of off-the-shelf planning techniques in LLMs (Valmeekam et al., 2023b). As demonstrated by (Valmeekam et al., 2023b), even in a seemingly simple common-sense domain like Blocksworld that humans usually find easy to solve, LLMs are evaluated to be quite ineffective in planning autonomously.
An interesting result shown by (Valmeekam et al., 2023b) is when taking the solution generated by LLMs, which is generally incorrect, as a seed plan to be repaired by an off-the-shelf planner, e.g., LPG (Gerevini & Serina, 2002), a significant improvement in search steps can be seen over the empty plan provided as a seed plan for the planner. This indicates that LLMs can indeed provide some helpful information (e.g., in some sense of heuristics) for planning, even though they cannot solve planning problems solely. Inspired by the result of loosely using plans generated by LLMs as seed plans, we are curious if it is possible to “dig” more helpful information from LLMs to assist planning deeply, e.g., by inserting LLMs into planning frameworks. By doing this, we aim to answer the question: what roles can be played exactly by LLMs in planning? Indeed, there have been attempts to explore off-the-shelf planning techniques to help LLMs solving planning problems (Liu et al., 2023). Similar to (Valmeekam et al., 2023b), they only view planners as black-boxes without deepening the integration of LLMs in planning frameworks.
To do this, we investigate the effectiveness of embedding LLMs into one of the well-known planning frameworks, graph-based planning (Blum & Furst, 1997). We propose a novel LLMs-based planning framework with LLMs embedded in two levels of the planning framework (namely LLMs4Plan), i.e., (a) selecting promising actions in each action-level to expand the planning graph, and (b) building non-mutual action sets in each action-level of the expanded planning graph, which correspond to two critical steps that influence the efficiency and effectiveness of graph planning.
(a) Expantion
(b) Backtracking
@@
For example, as shown in Figure 1(a), there could be many actions in “action-level 1” after expanding the planning graph based on “state-label 0” with Graphplan (Blum & Furst, 1997). We expect that LLMs will help select only a few promising actions in each action level, e.g., $\{a_{1},a_{2},a_{3}\}$ in “action-level 1”. In Figure 1(b), when backtracking from “state-level $K$” that contains goals, there could be many candidate sets of actions to be explored (e.g., “action set 1”, “action set 2”, “action set 3”) — actions in each candidate set are not mutual with each other. It is particularly time-consuming to search all of the valid candidate sets based on the mutual constraints in each action-level. We expect that LLMs are capable of helping select only a few candidate sets to be backtracked, e.g., only “action set 1” is selected to be backtracked by LLMs.
Specifically, in LLMs4Plan, for each action-level $i$, we first automatically generate prompts based on propositions in state-level $i-1$, goals and domain models when using LLMs to select actions to expand planning graphs in action-level $i$. After that, when backtracking from the last state-level of expanded planning graphs, i.e., the last state-level in expanded planning graphs contains goals, we automatically generate prompts based on propositions in state-level $j$ (from which backtracking is conducted), propositions from state-level 0, mutual constraints in action-level $j$, and domain models. We embed the above two components into one of the well-known off-the-shelf graph planners, Graphplan. We study the effectiveness of different cases of adding or removing the above one or two components in Graphplan to see the significance of roles LLMs play in the graph planning framework.
Through this study, we provide new clues for how to deeply embed LLMs into off-the-shelf planning frameworks, i.e., first identifying critical steps (generally time-consuming ones) in specific planning frameworks, and then designing proper prompt generation procedure to be embedded into the frameworks. We verify that soly relying on LLMs to do planning is far from a good option, while leveraging LLMs to help deal with some critical steps in the graph planning framework is possible.
@@ 2 Related work
LLMs as planners: Many works have investigated how to leverage LLMs to accomplish planning tasks, such as Chain-of-Thoughts (Wei et al., 2022), Tree-of-Thoughts (Yao et al., 2023a) and Zero-Shot Planner (Huang et al., 2022a), which utilize prompts to guide LLMs for generating action sequences for planning related tasks. Approaches such as HuggingGPT (Shen et al., 2023) and Chameleon (Lu et al., 2023) aim to generate initial plans using different tools and then call the corresponding APIs for execution by augmenting LLMs with plug-and-play modules for compositional reasoning. They designed an LLM-based planner to assemble a sequence of tools to execute to generate final solution plans for planning tasks. There are also some works that prompt LLMs to compose plans in the form of PDDL (Planning Domain Definition Language) in planning community (Liu et al., 2023; Pallagani et al., 2023). All of the above-mentioned approaches generally synthesize plans without consideration of feedback from external environments. In order to consider environmental feedback, SayCan (Ichter et al., 2022), ReAct (Yao et al., 2023b), Reflexion (Shinn et al., 2023) and Inner Monologue (Huang et al., 2022b) allow LLMs to take single-step actions according to the environmental feedback, assuming LLM-generated initial plans are correct without any adaptation of them. They only adjust the immediate action being executed and are prone to fall into local sub-optimal actions without considering long-term goals. In consideration of removing this assumption, the evaluation of planning capabilities in the literature involves the user incrementally interacting with LLMs, and re-prompting it to point out flaws in its plans, with the hope that the LLM eventually reaches an executable plan (Huang et al., 2022b; Yao et al., 2023b; Raman et al., 2022; Wang et al., 2023; Sun et al., 2023; Silver et al., 2023). Such evaluations are notoriously with the actual planning being done by humans in the loop rather than LLMs themselves. Instead of considering human-in-the-loop, (Valmeekam et al., 2023a) built another framework for automatically evaluating planning capabilities of LLMs by leveraging automated planning models and tools to generate the queries and validate answers from LLMs. All of the above-mentioned approaches view LLMs as a sole planner rather than being leveraged as components and embedded into off-the-shelf planning frameworks, which means the planning capability of planning frameworks is not leveraged by them.
LLMs finetuned to be Planners: Another issue of broad interest in LLMs is to investigate how to finetune LLMs to be capable of planning. Some works (Reed et al., 2022; Brohan et al., 2023) collected paired visual, language and action data and trained an enormous neural networks for tackling long-horizon tasks. Collecting those data, however, is often expensive and difficult to scale up. Different from training LLMs from scratch, finetuning off-the-shelf LLMs (Driess et al., 2023; Li et al., 2022) to deal with visual and language related tasks using task-specific robot demonstrations is an option to alleviate issues of data collection. It is, however, still problematic since finetuning most high-performing language models, such as GPT3.5/4 (Ouyang et al., 2022) and PaLM (Chowdhery et al., 2023), is currently impossible, as the model weights are not open-sourced. Besides, paired vision and language robotics demonstrations, which are still needed for finetuning, are not readily available in practice. Considering this, (Ajay et al., 2023) propose compositional foundation models for hierarchical planning, where a foundation model is a composition of different expert models trained on language, vision, and action data individually. Given an abstract language instruction describing the desired task, they use a large language model to generate a sequence of sub-tasks and then use a large video diffusion model to capture geometric and physical information about the world and generate a more detailed plan. Finetuning LLMs to be a planner indicates training data are required to be available and the planning capability of finetuned LLMs to be transferred to other tasks is still an open issue. Instead of finetuning LLMs, we aim to leverage both benefits of LLMs and planning frameworks to solve planning tasks without any training data.
@@
LLMs as components in planners: Some studies try to use LLMs as heuristics (Zhang et al., 2023) or transition function (Hao et al., 2023) in MCTS, boosting the performance in coding or small-scale reasoning. Utilizing LLMs as a world model in depth, however, is not discussed (Zhao et al., 2023). LLMs can also provide a commonsense model of the world in addition to a policy that acts on it. The world model and the policy are combined in Monte Carlo Tree Search (MCTS) to scale up task planning (Zhao et al., 2023), which is demonstrated to be effective by embedding LLMs in the MCTS framework. The idea is similar to our work in the sense that LLMs are embedded into an off-the-shelf framework in depth. However, they aim to learn MCTS policies from interactions with environments, while our work aims to solve planning problems with LLMs embedded into an off-the-shelf planner, planning-graph planner, without any interactions or learning from environments.
@@ 3 Problem Formulation
In this work we consider classical planning problems specified in the form of STRIPS language (Fikes & Nilsson, 1971). Similar idea can be extended into more expressive planning language such as PDDL (Ghallab et al., 1998). A STRIPS domain is composed of a set of action models $\mathcal{A}$, each of which is a quadruple $\langle a,\textrm{PRE}(a),\textrm{ADD}(a),\textrm{DEL}(a)\rangle$, where $a$ is an action name with zero or more parameters, $\textrm{PRE}(a)\subseteq\mathcal{L}$ is a precondition list indicating the conditions under which $a$ can be applied, $\textrm{ADD}(a)\subseteq\mathcal{L}$ is an adding list and $\textrm{DEL}(a)\subseteq\mathcal{L}$ is a deleting list indicating the effects of $a$. $\mathcal{L}$ is a set of atoms, each of which is composed of a predicate with zero or more parameters. We define a planning problem as $\mathcal{P}=\langle\mathcal{R},s_{0},g,\mathcal{A}\rangle$, where $s_{0}\subseteq\mathcal{R}$ is an initial state and $g\subseteq\mathcal{R}$ is a goal. $\mathcal{R}$ is a set of propositions. A solution plan $\pi=\langle A,C\rangle$ to $\mathcal{P}$ consists of a set of actions $A$ and a set of partial order constraints $C$ among actions. Some actions in $A$ are allowed to be executed in parallel so long as there are no conflicts among them, e.g., there are conflicts between two actions if one deletes a precondition or an adding effect of the other (more detailed description of conflicts will be given in Section 4.2). A partial order plan can be equivalently stretched into a total order linear plan, which means both of them can successfully transit initial state $s_{0}$ to goal $g$. An intuitive example of our planning problem is as shown below.
=========== Beginning of the example ============
Example 1: Suppose we would like to clean a bedroom using a vacuum which is placed in a tool room. We can formulate the problem $\mathcal{P}=\langle\mathcal{R},s_{0},g,\mathcal{A}\rangle$ in the form of planning language as shown below (note that we assume there is no parameter for each predicate and action for simplicity):
  • the set of propositions: =$\mathcal{R}=\{dirty(),toolroom(),clean(),bedroom()\}$, which indicates the set of all propositions in the problem.
  • initial state: $s_{0}$={dirty(), toolroom()}, which indicates the bedroom to be cleaned is dirty, and the tool vacuum is in the tool room.
  • goal: $g=\{clean(),toolroom()\}$, which indicates the bedroom is clean, and the tool vacuum is back to the tool room.
  • action models: $\mathcal{A}$ is as shown below:
@@
    $vacuum()$ $dirty()$, $clean()$,
    $bedroom()$$\neg dirty()$
    $move2tr()$ $bedroom()$ $toolroom()$,
    $\neg bedroom()$
    $move2br()$ $toolroom()$ $bedroom()$,
    $\neg toolroom()$
    Action vacuum aims to vacuuming the bedroom, the preconditions of which are the bedroom is dirty and the vacuum-cleaner is in the bedroom. The effects of vacuum are the bedroom is clean (deleting dirty() as well). Action $move2tr$ aims to move the vacuum-cleaner to the tool room, the precondition of which is the vacuum-cleaner is in the $bedroom$. The effects are the vacuum-cleaner is in the tool room and not in the bedroom. Similarly, action $move2br$ aims to move the vacuum-cleaner to the bedroom, the precondition of which is the vacuum-cleaner is in the tool room. The effects are the vacuum-cleaner is in the bedroom and not in the tool room.
The solution plan $\pi$ to the problem $\mathcal{P}$ is: $move2br(),vacuum(),move2tr()$.
============= End of the example ==============
@@ 4 Our LLMs4Plan approach
An overview of our LLMs4Plan approach is as shown in Algorithm 1. In Step 1, we first build a planning graph $PG$ with the help of LLMs given the planning problem $\mathcal{P}$ such that the goal is included by the last state-level in $PG$. If the number of levels in $PG$ is less than the maximal levels $MaxIter$, we build mutual constraints $\mathcal{C}$ and sort the candidate sets of actions, resulting in the list of candidate sets $\mathbb{A}$ — each of which satisfies $\mathcal{C}$ — with the help of LLMs. After that we make a deep first search based on $\mathbb{A}$ to get $\pi$. If $\pi$ is not Failure, it is returned as the solution plan; otherwise, we expand $PG$ with one more level using LLMs and repeat Steps 2-11 of the algorithm until the maximal level $MaxIter$ is reached or a solution plan $\pi$ is found and returned in Step 7. In the following subsections, we will detail main steps of Algorithm 1.
Algorithm 1 An overview of our LLMs4Plan approach
Input: Planning problem $\mathcal{P}$
@@
1:  Build graph: $(PG,k)=buildGraphLLMs(\mathcal{P})$
2:  while $k<MaxIter$ do
3:     Build constraints: $\mathcal{C}=buildConstraints(PG)$
4:     Sort action sets: $\mathbb{A}=sortActionsLLMs(PG,\mathcal{C})$
5:     Search solutions: $\pi=depthFirstSearch(\mathbb{A},PG)$
6:     if $\pi\neq$ Failure then
7:        return $\pi$
8:     end if
9:     Expand graph: $PG\leftarrow expandGraphLLMs(PG)$
10:     $k\leftarrow k+1$
11:  end while
12:  return Failure

### v2

@@ 1 Introduction
Plan synthesis aims to generate a course of actions or policies to transit given initial states to goal states, provided domain models that could be designed by experts or learnt from training data [1] or interactions with the world [15, 14]. It is a time- and space-consuming open issue in the planning community [6]. Intrigued by the claims of emergent planning capabilities in large language models (LLMs), works have been proposed to investigate the planning effectiveness of LLMs, without considering any utilization of off-the-shelf planning techniques in LLMs [27]. As demonstrated by [27], even in a seemingly simple common-sense domain like Blocksworld that humans usually find easy to solve, LLMs are evaluated to be quite ineffective in planning autonomously.
An interesting result shown by [27] is when taking the solution generated by LLMs, which is incorrect, as a seed plan to be repaired by an off-the-shelf planner, e.g., LPG [4], a significant improvement in search steps can be attained over the result when an empty plan provided as a seed plan for the planner. This indicates that LLMs can indeed provide some helpful information (e.g., in some sense of heuristics) for planning, even though they cannot solve planning problems solely. Inspired by the result of loosely using plans generated by LLMs as seed plans, we are curious if it is possible to “dig” more helpful information from LLMs to assist planning deeply, e.g., by inserting LLMs into planning frameworks. By doing this, we aim to answer the question: what roles can be played exactly by LLMs in planning? Indeed, there have been attempts to explore off-the-shelf planning techniques to help LLMs solving planning problems [17]. Similar to [27], they only view planners as black-boxes without deepening the integration of LLMs in planning frameworks.
To do this, we investigate the effectiveness of embedding LLMs into one of the well-known planning frameworks, graph-based planning [2]. We propose a novel LLMs-based planning framework with LLMs embedded in two phases of the planning framework (namely LLMs4Plan). The first phase is to propose promising actions in “action-levels” of the planning graph using LLMs. The second phase is to propose non-mutual action sets using LLMs when backtracking the planning graph. Note that the two phases correspond to two critical steps that influence the efficiency and effectiveness in graph planning.
@@
For example, as shown in Figure 1(a), there could be a large number of actions in “action-level 1” when expanding “state-label 0” with Graphplan [2]. We aim to exploit LLMs to help select a small subset of promising actions, e.g., $\{a_{1},a_{2},a_{3}\}$ are selected in Figure 1(a). In Figure 1(b), when backtracking from “state-level $K$” that includes goals, there could be a large number of candidate sets of actions to be explored (e.g., “action set 1”, “action set 2”, “action set 3”) — actions in each candidate set are not mutually exclusive with each other (two actions are mutually exclusive if they are not allowed to be executed at the same time, e.g., actions “pick up object A” and “put down object A” are mutually exclusive). It is particularly time-consuming to search all of the valid candidate sets based on mutual constraints in each action-level. We expect LLMs are capable of selecting a small number of candidate sets to be backtracked, e.g., only “action set 1” is selected to be backtracked by LLMs as shown in Figure 1(b).
Specifically, in LLMs4Plan, for each action-level $i$, we first automatically generate prompts based on propositions in state-level $i-1$, goals and domain models, and feed the prompts to LLMs to select actions for generating action-level $i$. After that, when backtracking from the last state-level $K$ in the expanded planning graph, which includes the goal, we automatically generate prompts based on propositions in state-level $K$, propositions in the initial state (i.e., state-level 0), mutual constraints in action-level $K$, and domain models, and feed the prompts to LLMs to select action sets for backtracking. We embed the above two components into one of the well-known off-the-shelf graph planners, Graphplan [2]. We study the effectiveness of different cases of adding or removing the above one or two components in Graphplan to see the significance of roles LLMs play in the graph planning framework.
Through this study, we provide new clues for how to deeply embed LLMs into off-the-shelf planning frameworks, i.e., first identifying critical steps (generally time-consuming ones) in specific planning frameworks, and then designing proper prompt generation to be embedded into the frameworks. We verify that soly relying on LLMs to do planning is far from a good option, while leveraging LLMs to help deal with some critical steps in the graph planning framework is possible.
@@ 2 Related work
LLMs as planners: There have beem works leveraging LLMs to assisting planning tasks, such as Chain-of-Thoughts [29], Tree-of-Thoughts [30] and Zero-Shot Planner [10], which utilize prompts to guide LLMs for generating action sequences for planning related tasks. Approaches such as HuggingGPT [21] and Chameleon [18] aim to generate initial plans using different tools and then call the corresponding APIs for execution by augmenting LLMs with plug-and-play modules for compositional reasoning. They designed an LLM-based planner to assemble a sequence of tools to execute to generate final solution plans for planning tasks. There are also some works that prompt LLMs to compose plans in the form of PDDL (Planning Domain Definition Language) in planning community [17, 19]. All of the above-mentioned approaches generally synthesize plans without consideration of feedback from external environments. In order to consider environmental feedback, SayCan [12], ReAct [31], Reflexion [22] and Inner Monologue [11] allow LLMs to take single-step actions according to the environmental feedback, assuming LLM-generated initial plans are correct without any adaptation of them. They only adjust the immediate action being executed and are prone to fall into local sub-optimal actions without considering long-term goals. In consideration of removing this assumption, the evaluation of planning capabilities in the literature involves the user incrementally interacting with LLMs, and re-prompting it to point out flaws in its plans, with the hope that the LLM eventually reaches an executable plan [11, 31, 20, 28, 25, 24]. Such evaluations are notoriously with the actual planning being done by humans in the loop rather than LLMs themselves. Instead of considering human-in-the-loop, [27] built another framework for automatically evaluating planning capabilities of LLMs by leveraging automated planning models and tools to generate the queries and validate answers from LLMs. All of the above-mentioned approaches view LLMs as a sole planner rather than being leveraged as components and embedded into off-the-shelf planning frameworks, which means the planning capability of planning frameworks is not leveraged by them.
@@
LLMs as components in planners: There are also works using LLMs as heuristics [32] or transition function [8] in MCTS, boosting the performance in coding or small-scale reasoning. LLMs can also provide a commonsense model of the world in addition to a policy that acts on it. The world model and the policy can be combined into Monte Carlo Tree Search (MCTS) to scale up task planning [33], which is demonstrated to be effective by embedding LLMs in the MCTS framework. The idea is similar to our work in the sense that LLMs are embedded into an off-the-shelf framework in depth. However, they aim to learn MCTS policies from interactions with environments, while our work aims to solve planning problems with LLMs embedded into an off-the-shelf planner, planning-graph planner, without any interactions or learning from environments. [7] investigated effectiveness of the generated PDDL action models from LLMs for downstream planning tasks. Likewise, it is different from our work in the sense that we assume the PDDL action models are already known and focus on exploring LLMs to speed up the planning efficiency.
@@ 3 Problem Formulation
In this work we consider classical planning problems specified in the form of STRIPS [3]. Similar ideas can be extended into more expressive planning language such as PDDL [5]. Let $\mathcal{L}$ be a set of atoms, each of which is composed of a predicate with zero or more parameters (e.g., clean(room) is an atom indicating room is clean). A STRIPS domain is composed of a set of action models $\mathcal{A}$, each of which is a quadruple $\langle a,\textrm{PRE}(a),\textrm{ADD}(a),\textrm{DEL}(a)\rangle$, where $a$ is an action name with zero or more parameters, $\textrm{PRE}(a)\subseteq\mathcal{L}$ is a precondition list indicating the conditions under which $a$ can be applied, $\textrm{ADD}(a)\subseteq\mathcal{L}$ is an adding list and $\textrm{DEL}(a)\subseteq\mathcal{L}$ is a deleting list indicating the effects of $a$. Let $\mathcal{R}$ be a set of propositions, which are instances of atoms in $\mathcal{L}$. We define a planning problem as $\mathcal{P}=\langle\mathcal{R},s_{0},g,\mathcal{A}\rangle$, where $s_{0}\subseteq\mathcal{R}$ is an initial state and $g\subseteq\mathcal{R}$ is a goal. A solution $\pi$ to the planning problem is a sequence of actions that transit initial state $s_{0}$ to goal $g$. An intuitive example of our planning problem is as shown below.
Suppose we would like to clean a bedroom using a vacuum which is placed in a tool room. We can formulate the problem $\mathcal{P}=\langle\mathcal{R},s_{0},g,\mathcal{A}\rangle$ in the form of STRIPS (note that we assume there is no parameter for each predicate and action for simplicity since there is only one tool, one bedroom and one toolroom). The set of propositions $\mathcal{R}$ is represented by =$\mathcal{R}=\{dirty(),toolroom(),clean(),bedroom()\}$. Initial state $s_{0}$ is represented by $s_{0}$ = {dirty(), toolroom()}, which indicates the “bedroom” is dirty, and the tool “vacuum” is in the tool room (i.e., “toolroom”). The goal $g$ is represented by $g=\{clean(),toolroom()\}$, which indicates the “bedroom” is clean, and the tool “vacuum” is back to the tool room. The set of action models $\mathcal{A}$ is represented as follows:
@@
$vacuum()$ $dirty()$, $bedroom()$ $clean()$, $\neg dirty()$
$move2tr()$ $bedroom()$ $toolroom()$, $\neg bedroom()$
$move2br()$ $toolroom()$ $bedroom()$, $\neg toolroom()$
Action $vacuum()$ aims to vacuuming “bedroom”, the preconditions of which are “bedroom” is dirty and the vacuum-cleaner is in “bedroom”. The effects of $vacuum()$ are adding $clean()$ to the state where $vacuum()$ is executed, indicating “bedroom” is clean, and deleting $dirty()$ (i.e., $\neg dirty()$) from the state, indicating “bedroom” is not dirty anymore. Action $move2tr()$ aims to move the vacuum-cleaner to “toolroom”, the precondition of which is $bedroom()$ indicating the vacuum-cleaner is in “bedroom”. The effects are adding $toolroom()$ indicating the vacuum-cleaner is in “toolroom”, and deleting $bedroom()$, indicating the vacuum-cleaner is not in “bedroom”. Similarly, action $move2br$ aims to move the vacuum-cleaner to “bedroom”, the precondition of which is $toolroom()$, indicating the vacuum-cleaner is in “toolroom”. The effects are the vacuum-cleaner is adding $bedroom()$, indicating the vacuum-cleaner is in “bedroom”, deleting $toolroom()$, indicating the vacuum-cleaner is not in “toolroom”. A solution $\pi$ to the problem $\mathcal{P}$ is $move2br(),vacuum(),move2tr()$.
@@ 4 Our LLMs4Plan approach
An overview of our LLMs4Plan approach is shown in Algorithm 1. In Step 3, the pruning possibility $\kappa_{i}$ is decreased as the exponent $i$ increasing. In Step 5, we expand planning graph $PG^{r}$ with one more level using LLMs to prune actions based on pruning possibility $\kappa_{i}$ and planning problem $\mathcal{P}$. In Steps 7, if goal $g$ is not included by the last state-level in $PG^{r}$, i.e., $Satisfied(g,PG^{r})$ is false, we continue to Step 4. In Step 8, we build a set of mutual constraints $\mathcal{C}$ based on $PG^{r}$, i.e., $buildConstraints(PG^{r})$. In Step 9, we sort sets of actions based on constraints $\mathcal{C}$ using LLMs, i.e., $sortActionsLLMs(PG^{r},\mathcal{C}$. In Step 10, we search solution $\pi$ based on the sorted action sets $\mathbb{A}$ using depth-first search. In the following subsections, we will address our LLMs4Plan in detail.
Algorithm 1 An overview of our LLMs4Plan
Input: Planning problem $\mathcal{P}$, pruning possibility $\kappa_{0}$
@@
1:  $PG^{r}=\emptyset$
2:  for $i=1$ to $N$ do
3:     $\kappa_{i}=(\kappa_{0})^{i}$, $k=1$
4:     while $k<K$ do
5:        $PG^{r}\leftarrow expandGraphLLMs(PG^{r},\mathcal{P},\kappa_{i})$
6:        $k=k+1$
7:        if $Satisfied(g,PG^{r})=false$, then continue
8:        $\mathcal{C}=buildConstraints(PG^{r})$
9:        $\mathbb{A}=sortActionsLLMs(PG^{r},\mathcal{C})$
10:        $\pi=depthFirstSearch(\mathbb{A},PG^{r})$
11:        if $\pi\neq$ Failure, then return $\pi$
12:     end while
13:  end for
14:  return Failure